OpenStudy (anonymous):

Will Fan And Medal. Looking For A Straight Answer. What is the value of the expression 22 + 82 ÷ 22? 8 10 17 20

OpenStudy (anonymous):

\[22+\frac{82}{22}=\frac{283}{11} \]

OpenStudy (anonymous):

@robtobey i dont get it.. anyway you can give me a straight answer?

OpenStudy (anonymous):

Division is calculated before addition. 283/11 is correct.

OpenStudy (anonymous):

Divide 82 by 22. This equals 3.7 Add 3.7 to 22 So 3.7+22= 25.7 your answer is 25.7

OpenStudy (anonymous):

@sophiamariess098 25.7 isnt one of the answer choices.

OpenStudy (anonymous):

@robtobey im sorry but im looking more for a straight answer..

OpenStudy (anonymous):

I presented the answer in the form of a fraction. @sophiamariess098 posted the answer as a decimal number. Perhaps 22 + 82 ÷ 22 is not the intended problem expression.

OpenStudy (anonymous):

what are the choises?